A solution contains 0.71 mol of aluminum chloride in a 910 mL solution. Calculate the molarity of all the ions (Al and Cl) in the solution.



1
Expert's answer
2021-08-02T12:27:09-0400

Solution:

Convert mL to L:

1 L = 1000 mL

Thus, (910 mL solution) × (1 L / 1000 mL) = 0.91 L solution


One formula unit of aluminum chloride (AlCl3) contains one aluminum cation (Al3+) and three chloride anions (Cl).

This means that when 1 mol of aluminum chloride is dissolved in water, 1 mol of aqueous aluminum ions and 3 mol of aqueous chloride ions are formed.

AlCl3(aq) → Al3+(aq) + 3Cl(aq)

From this it follows that the number of moles of Al3+ and Cl- present in this solution will be:

Moles of Al3+ = Moles of AlCl3 = 0.71 mol Al3+

Moles of Cl= 3 × Moles of AlCl3 = 3 × (0.71 mol) = 2.13 mol Cl


Molarity = Moles / Solution volume

Therefore,

Molarity of Al3+ = (0.71 mol Al3+) / (0.91 L) = 0.78 mol/L Al3+ = 0.78 M Al3+

Molarity of Cl= (2.13 mol Cl) / (0.91 L) = 2.34 mol/L Cl = 2.34 M Cl

Molarity of AlCl3 = (0.71 mol AlCl3) / (0.91 L) = 0.78 mol/L AlCl3 = 0.78 M AlCl3


Answer:

Molarity of Al3+ is 0.78 M

Molarity of Cl is 2.34 M